A sample of gas has a volume of 1.73L at a pressure of 860 mmHg. What must the pressure be, in mmHg, on this sample for the volume to change to 2.40L?


1
Expert's answer
2021-01-05T06:02:00-0500

Аccording the gas Law p1*V1=p2V2

From which p2=p1*V1/V2= 1.73 *860/2.4= 619.9 mmHg